$SPAD/src/input frame.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool frame.output
)set message test on
)set message auto off
)clear all
--S 1 of 24
)frame new testframe
--R
--E 1
--S 2 of 24
)frame names
--R
--R   The names of the existing frames are:
--R            testframe
--R            initial
--R      The current frame is the first one listed.
--E 2
--S 3 of 24
)frame next
--R
--E 3
--S 4 of 24
)frame names
--R
--R   The names of the existing frames are:
--R            initial
--R            testframe
--R      The current frame is the first one listed.
--E 4
--S 5 of 24
)frame next
--R
--E 5
--S 6 of 24

)frame names
--R
--R   The names of the existing frames are:
--R            testframe
--R            initial
--R      The current frame is the first one listed.
--E 6
--S 7 of 24
)frame next
--R
--E 7
--S 8 of 24
)frame drop
--R
--E 8
--S 9 of 24
)frame names
--R
--R   The names of the existing frames are:
--R            testframe
--R      The current frame is the first one listed.
--E 9
--S 10 of 24
)frame new testframe2
--R
--E 10
--S 11 of 24
a:=1
--R
--R
--R   (1)  1
--R                                                        Type: PositiveInteger
--E 11
--S 12 of 24
)frame names
--R
--R   The names of the existing frames are:
--R            testframe2
--R            testframe
--R      The current frame is the first one listed.
--E 12
--S 13 of 24
)frame next

--R
--E 13
--S 14 of 24
a
--R
--R
--R   (1)  a
--R                                                            Type: Variable(a)
--E 14
--S 15 of 24
)frame import testframe2 a
--R
--R   Import from frame testframe2 is complete. Please issue )display all
--R      if you wish to see the contents of the current frame.
--E 15
--S 16 of 24
)display all
--R
--RProperties of % :
--R   Value (has type Variable(a)):  a
--RProperties of %e :
--R   This is a system-defined macro.
--R   macro %e () == exp(1)
--RProperties of %i :
--R   This is a system-defined macro.
--R   macro %i () == complex(0,1)
--RProperties of %infinity :
--R   This is a system-defined macro.
--R   macro %infinity () == infinity()
--RProperties of %minusInfinity :
--R   This is a system-defined macro.
--R   macro %minusInfinity () == minusInfinity()
--RProperties of %pi :
--R   This is a system-defined macro.
--R   macro %pi () == pi()
--RProperties of %plusInfinity :
--R   This is a system-defined macro.
--R   macro %plusInfinity () == plusInfinity()
--RProperties of SF :
--R   This is a system-defined macro.
--R   macro SF () == DoubleFloat()
--RProperties of a :
--R   Value (has type PositiveInteger):  1
--E 16
--S 17 of 24
a

--R
--R
--R   (2)  1
--R                                                        Type: PositiveInteger
--E 17
--S 18 of 24
)set message prompt frame
--R
--E 18
--S 19 of 24
)frame next
--R
--E 19
--S 20 of 24
)frame next
--R
--E 20
--S 21 of 24
)frame next
--R
--E 21
--S 22 of 24
)frame names
--R
--R   The names of the existing frames are:
--R            testframe2
--R            testframe
--R      The current frame is the first one listed.
--E 22
--S 23 of 24
)frame drop testframe2
--R
--E 23
--S 24 of 24
)frame names
--R
--R   The names of the existing frames are:
--R            testframe
--R      The current frame is the first one listed.
--E 24
)spool
)lisp (bye)

———- 7

References^{[1] nothing}

8